Assume that a sample is used to estimate a population mean μμ. Find the 98% confidence interval for a sample of size 40 with a mean of 72.3 and a standard deviation of 14.5. Enter your answer as an open-interval (i.e., parentheses) accurate to one decimal place (because the sample statistics are reported accurate to one decimal place).

98% C.I. =

Answer should be obtained without any preliminary rounding. However, the critical value may be rounded to 3 decimal places.
Answer #1

Solution :

degrees of freedom = n - 1 = 40 - 1 = 39

t/2,df = 2.426

Margin of error = E = t/2,df * (s /n)

= 2.426* ( 14.5 / 40)

Margin of error = E = 5.6

The 98% confidence interval estimate of the population mean is,

  ± E

72.3 ± 5.6

(66.7 , 77.9)